OpenStudy (anonymous):

What is the slope and y intercept of y=-1/3x+3

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = mx + b}\) Slope = m y-intercept = b

Parth (parthkohli):

What's m here?

OpenStudy (anonymous):

-1/3

Parth (parthkohli):

yeah, and what's b?

OpenStudy (anonymous):

3?

Parth (parthkohli):

Yep

Parth (parthkohli):

So slope = -1/3 and y-intercept = 3

OpenStudy (anonymous):

Ohh I feel so dumb right now

Parth (parthkohli):

Why? You already answered it :)